1/1

William Wordsworth,  William Wordsworth,
Die Narzissen  The Daffodils

tradukita de Hans Hennecke

Argiope-informo: La aspekto de tiu ĉi versio estas barita. Supozeble la aŭtoro decidis tion.

I wander’d lonely as a cloud That floats on high o’er vales and hills, When all at once I saw a crowd, A host of golden daffodils, Beside the lake, beneath the trees Fluttering and dancing in the breeze.

Continuous as the stars that shine And twinkle on the milky way, They stretch’d in never-ending line Along the margin of a bay: Ten thousand saw I at a glance Tossing their heads in sprightly dance.

The waves beside them danced, but they Out-did the sparkling waves in glee: - A poet could not but be gay In such a jocund company: I gazed - and gazed - but little thought What wealth the show to me had brought.

For oft, when on my couch I lie In vacant or in pensive mood, They flash upon that inward eye Which is the bliss of solitude; And then my heart with pleasure fills And dances with the daffodils.

Traduko de la Angla poemo “The Daffodils” de William Wordsworth (∗1770 – †1850) en la Germanan de Hans Hennecke.

Arg-39-75 (2004-02-01 18:36:59)

Germanigo de la origina anglalingva teksto de WILLIAM WORDSWORTH, fare de Heneko, Johano (HANS HENNECKE). (Mi trovis tiun ci poemon en la libro �Snapshots of Britain“ / �Britannien in kleinen Geschichten“, rakontita de JOY BROWNING, eldonejo Deutscher Taschenbuchverlag, ISBN 3-423-09319-6.)

Verkinto de tiu ĉi Angla poemo estas William Wordsworth (∗1770 – †1850).

Arg-39-73 (2003-10-17 22:02:21)

Origina anglalingva teksto de WILLIAM WORDS-WORTH (1770 - 1850)

http://www.poezio.net/version?poem-id=39&version-id=75,73